OpenStudy (anonymous):

Find the vertical and horizontal asymptotes. y=(2x-5)/(|3x+2|)

OpenStudy (cwrw238):

the vertical asymptote is where the denominator is zero

OpenStudy (cwrw238):

solve 3x + 2 = 0 to find vertical asymptote

OpenStudy (anonymous):

using the limit :horizontal asymptotes.:\[\lim_{x \rightarrow \infty}(2x-5)/(|3x+2|)=\frac{ 2 }{3 }\],y=2/3 vertical asymptote:3x+2=0,x=-2/3

OpenStudy (anonymous):

Thank you :)